OpenStudy (anonymous):

What is x^2+7x+12/ 3x+9 reduced to its simpelest form?

OpenStudy (anonymous):

x^2+4x+3x+12=(x+4)(x+3) so the expression is (x+4)(x+3)/(3(x+3)) =(x+4)/3